# Comparative Study of State-based Neural Networks for Virtual Analog Audio Effects Modeling

### v5

@@ 3 Methods
This study aims to identify which type of recurrent layer or architectural configuration accurately models different types of audio effects, if any. These insights can be used to understand the goodness and weakness of these models and determine which can better suit each audio effect. The states-based architectures offer advantages due to their ability to capture time dependencies in the data by utilizing internal states rather than solely relying on the input received at each iteration. In addition, we restrict the experimentation to relatively small networks fed by a small number of input segments, which present low computational complexity and low input-output latency, allowing implementation on real-time consumer-grade digital audio systems. The designed models are applied to different audio effects and finally evaluated using different metrics. To be noticed, the models presented here, having limited parameters, are not expected to lead to perfectly accurate emulation but a good starting point.
In the rest of this section, we first detail the four architectures used in this comparative study; then, we present the metrics we used for evaluating the trained models and the specific audio effects from which we collected the datasets. Finally, explain the process we used to train and test the models.
@@ 3.1 Architectures
The following architectures, based on different types of recurrent layers, are compared in this study: LSTM (a), ED (b), LRU (c), and S4D (d). The four architectures are illustrated in Figure 1. ED is based on the concept of internal state sharing  (Simionato & Fasciani,, 2022). These architectures follow the same design, consisting of a linear, fully connected layer, a recurrent layer, and a conditioning block.
The number of input samples is identical for all architectures and set to $64$. This choice represents a trade-off: on the one hand, there is the amount of past input information used to generate the current output, where more is typically better; on the other hand, there is the resulting input-output latency, which increases with the number of input samples. Our goal is to keep this latency below $1.33$ ms (i.e., $64$ samples at $48$ kH.). The input of all architectures is an array, including the $64$ most recent input sample, used to generate the single current output sample. This implies that a new input sample enters the buffer with each inference cycle, while the oldest is discarded. Therefore, the set of the $64$ most recent input samples constitutes the network input, which, after initial compression by the FC layer, is used to update the states in the recurrent layers. This approach aids the network in making predictions while maintaining a stateful design using truncated backpropagation through time. Predicting one sample at a time helps minimize the audible artifacts commonly associated with machine-learning models of audio effects, which arise from slight amplitude mismatches at the boundaries of consecutive output segments. Consequently, the model must perform predictions at an audio rate, one for each output audio sample. The models map a vector of $64$ sample in one scalar output sample, give the internal states $\boldsymbol{h}$:
@@
where $\boldsymbol{u_{n}}$ is the linear projected vector having a size equal to the number of hidden units of the fully connected layer. $\boldsymbol{W_{l}}$ and is the weights’ matrices, and $\boldsymbol{b_{l}}$ the vector bias terms.
Figure 1: The four compared architectures: LSTM, ED, LRU, S4D, and S6. All architectures present the same structure: a fully connected layer, a recurrent layer, and a conditioning layer. For the case of ED, the input vector is equally split into two vectors, $\boldsymbol{x_{e}}$ and $\boldsymbol{x_{d}}$.
The first two architectures use LSTMs as the recurrent layer, designed to improve the memory capacity of the vanilla RNN by incorporating internal mechanisms known as gates, regulating the flow of information. The LSTM architecture is governed by the following equations:
@@
where $\boldsymbol{h_{n}}$ is the hidden state vector, $\boldsymbol{f}$ is the forget gate, $\boldsymbol{i}$ the input gate, and $\boldsymbol{o}$ the output gate, and $\boldsymbol{c}$ is the cell state vector. The forget gate determines what information from previous steps should be kept or discarded. The input gate decides which information from the current step is relevant to add. The output gate determines the value of the next hidden state. Additionally, the cell state is updated by multiplying the forget vector and the input vector with new values that the neural network deems relevant. The forget vector and input vector are obtained using the sigmoid function, denoted by $\gamma$, which determines how much information should be kept. Values of 00 in this vector indicate information to be discarded, while a value of $1$ indicates information to be kept. Furthermore, the $\phi$ represents the hyperbolic tangent function, used to keep the value of the cell state between $-1$ and $1$. Finally, $\boldsymbol{W}$ and $\boldsymbol{U}$ are the weights’ matrices, and $\boldsymbol{b}$ the vector bias terms. After the LSTM layer, a linear FC layer reduces the dimensionality down to $4$, ensuring consistency with all other models before proceeding to the conditioning block.
In the case of ED, the input split into two equal-size parts, $\boldsymbol{x_{d}}$ containing $[x_{n},...,x_{n-31}]$ and $\boldsymbol{x_{e}}$ containing containing $[x_{n-32},...,x_{n-63}]$. The first part feds two separate convolutional layers, which compute the states, $[\boldsymbol{h},\boldsymbol{c}]$. These states are combined with the internal states of the LSTM layer computed during the previous iteration. A sigmoid function is applied to the LSTM internal states and then element-wise multiplied to $[h,c]$. This operation determines how much information should be taken from the computed $[h,c]$. The result of this operation represents the new LSTM’s internal states. This process is detailed in Figure 2.
Figure 2: The states sharing process in ED: the encoder input $\boldsymbol{x_{e}}$ is passed to two convolutional layers, which produce the states $[\boldsymbol{h},\boldsymbol{c}]$. These states are combined with $[\boldsymbol{h_{n-1}},\boldsymbol{c_{n-1}}]$, which are the LSTM states of the previous iteration. The sigmoid function is applied to $[\boldsymbol{h_{n-1}},\boldsymbol{c_{n-1}}]$ and the results are element-wise multiplied to $[\boldsymbol{h},\boldsymbol{c}]$. The outcome will be the new states for the LSTM layer $[\boldsymbol{h_{n}},\boldsymbol{c_{n}}]$.
The second part of the input vector is linearly transformed and fed into the second LSTM, following the same procedure as the other models.
The third architecture is based on LRU. Compared to a vanilla RNN, the LRU uses a linear activation function to compute the hidden state instead of a nonlinear function. This modification allows the memory capability to improve significantly. The LRU equations are the following:
@@
where $\boldsymbol{h_{n}}$ is the hidden state at time $n$ and acts as ’memory’ of the network, $\boldsymbol{o_{n}}$ the output vector. LRU keeps the recurrent layer linear but moves the nonlinearity to a subsequent fully connected layer. Additionally, to further enhance stability, the LRU incorporates a complex-valued diagonal recurrent matrix initialization and an exponential parameterization. At the same time, a normalization scheme for the hidden states is used to boost the efficiency and accuracy (Orvieto et al.,, 2023).
Lastly, the SSMs operate by considering a continuous-time representation of the state-space formulation that is discretized to obtain the following equations:
@@
In our study, we employ the diagonal state space model variant S4D  (Gu et al.,, 2022a), and newer S6 variant (Gu & Dao,, 2023).
The remarkable ability of S4D to capture long-range dependencies stems from their utilization of a specific state matrix $\boldsymbol{\bar{W_{A}}}$ known as the "HiPPO matrix" (Gu et al.,, 2020). This matrix is designed to encode all the past input history in $\boldsymbol{h_{n}}$, finding a map from $\boldsymbol{u_{n}}$ to a higher dimensional space $\boldsymbol{h_{n}}$ that represents the compression of the history. The matrix allows the model to be conceptualized as a convolutional model that decomposes an input signal onto an orthogonal system of smooth basis functions. In this way, the state $\boldsymbol{h_{n}}$ encodes the history of the signal. In the diagonal form, the S4D model is defined by parameterizing its state matrix as a diagonal matrix, boosting the model’s efficiency. By doing so, the basis kernels have closed-form formulas represented by normalized Legendre polynomials $L_{n}(t)$, resulting in the SSM decomposing the input signal $\boldsymbol{u_{n}}$ onto an infinitely long set of basis functions that are orthogonal with respect to an exponentially decaying measure.
In contrast, the S6 layer introduces input-dependent matrices. Specifically, $\boldsymbol{\bar{W_{B}}}$ and $\boldsymbol{\bar{W_{C}}}$ become dependent on the input and computed using a linear FC layer, while $\boldsymbol{\bar{W_{A}}}$ and $\boldsymbol{\bar{W_{D}}}$’s parameters are independent of the input and learned during the training process.
The recurrent layer’s output size will equal the number of hidden units of the layer. In the case of LRU and S4D, a fully connected layer with a hyperbolic tangent activation function also follows the recurrent layer to compensate for a lack of nonlinear activation functions in these layers:
$\boldsymbol{\hat{o}_{n}}=tanh(\boldsymbol{W_{nl}}\boldsymbol{o_{n}}+\boldsymbol{b_{nl}})$ (6)
The conditioning block is then applied just before the output layer, which consists of a linear fully connected layer with one unit, which is described in the following subsection. Placing the conditioning block after the recurrent layers rather than before has been found beneficial  (Fasciani et al.,, 2024a). This suggests that it is more beneficial for the networks to use the information given by the control parameters to project the output of the recurrent layer and determine the extent to which this information influences the final output rather than influencing the inference of the recurrent layer based on the control parameters.
@@
The number of units is chosen to have approximately $800$ trainable parameters among the architectures and effects since an identical number is achievable due to major differences in the layers. As a result, the selected LSTM and ED present $8$ units in the recurrent layer and $4$ in fully connected ones. LRU and S4D present $12$ units in the recurrent layer, and $6$ in fully connected ones. The number of trainable parameters also depends on the dimensional of the conditioning input $\boldsymbol{p}$, equal to the number of variable parameters in the datasets.
@@ 3.1.1 Conditioning Block
The conditioning block consists of the FiLM method and the GLU layer, as in (Fasciani et al.,, 2024a), and is illustrated in Figure 3.
@@
FiLM is a technique that applies an affine transformation to a vector based on conditioning values. The parameters vector $\boldsymbol{p}$, if applicable, is the input to a linear fully connected layer. The output of this layer is then split into two vectors of equal size: $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$. If $\boldsymbol{o_{n}}$ is the vector to which the transformation is to be applied, the output of the linear fully connected layer will have a length double of $\boldsymbol{o_{n}}$. Consequently, FiLM performs am affine transform on $\boldsymbol{o_{n}}$ using $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$.
Following the projection, it is common to apply a function, such as ReLU or sigmoid, to determine the amount of information that should be passed. In our approach, we replace the activation function with a GLU layer. This layer is more stable than ReLU and exhibits faster learning compared to sigmoid. Similarly to the previous step, the GLU layer consists of a linear fully connected layer that takes the FiLM output vector as input and computes a vector with twice its length. The resulting output is then split equally into two vectors: $\boldsymbol{q_{1}}$ and $\boldsymbol{q_{2}}$. A function is applied to $\boldsymbol{q_{2}}$, and the resulting output is multiplied element-wise with $\boldsymbol{q_{1}}$. Hence, the conditioning block is described by the following:
$\boldsymbol{o^{c}_{n}}=\boldsymbol{q_{1}}\odot softsign(\boldsymbol{q_{2}}),\text{ with }\boldsymbol{q_{1}}$ $=\boldsymbol{\dot{o}_{n}}\boldsymbol{K_{1}}+\boldsymbol{d_{1}}$
$\boldsymbol{q_{2}}$ $=\boldsymbol{\dot{o}_{n}}\boldsymbol{K_{2}}+\boldsymbol{d_{2}}$
$\boldsymbol{\dot{o}_{n}}$ $=(\boldsymbol{\theta}\boldsymbol{o_{n}}+\boldsymbol{\eta}).$ (8)
The GLU layer determines the flow of information through the network, acting as a logical gate. Typically, a sigmoid activation function is used for this purpose. However, in our approach, we introduce a softsign function instead. The softsign function controls the extent to which the control parameters should positively or negatively influence the final output. A value of 00 indicates a bypass, meaning no influence on the output. This allows for more flexibility.
@@ 3.2 Losses and Metrics
The models in this study take raw audio samples as input and produce raw audio samples as output. The most common losses used in this case include the Mean Squared Error (MSE), the Mean Absolute Error (MAE), and the Error-to-Signal Ratio (ESR). Based on our knowledge and supported by empirical experiments, there is no evidence suggesting that more complex loss functions could benefit the learning process, especially when modeling the transformation applied to a signal by a generic audio effect. While using specific and different loss functions for different effects may potentially yield better results, we decided to carry out this study using a fixed loss function across architectures and audio effects. This allows for a fair comparison between the trained models, as they all attempt to minimize the same loss function. Therefore, we selected the MSE for this comparative study as it helps to capture the dynamic aspects of the output signal more effectively  (Simionato & Fasciani,, 2023).
@@
In addition, our quantitative evaluation is based on various metrics to assess the models from different perspectives. In the following expressions, $y_{n}$ and $\hat{y}_{n}$ represent the target and predicted signal and the symbol $N$ represents the overall number of samples in the output signal.
@@
The first metric is the root-mean-square energy (RMSE) error. RMSE error provides insights into the energy deviation between the target and prediction. The losses and metrics presented so far calculate the average difference between the target and prediction, which may hide large errors that occur only during signal transients. Transients, such as sharp attacks or decays, may represent a small fraction of the overall duration of an audio signal. In this case, we may obtain a relatively low error even with all transients poorly modeled. Therefore, we have included a metric related to the spectral flux, which considers the differences between pairs of spectrums computed on consecutive overlapping windows, as used in the library  [^1] https://github.com/magenta/ddsp developed after  (Engel et al.,, 2020):
$\begin{split}M_{SF}=\Big{|}\Big{|}\Big{(}\Big{|}|STFT(y_{n})|-|STFT(y_{n-1})|\Big{|}\Big{)}\\ -\Big{(}\Big{|}|STFT(\hat{y}_{n})|-|STFT(\hat{y}_{n-1})|\Big{|}\Big{)}\Big{|}\Big{|}_{1}\end{split}$ (9)
where the Short-Time Fourier Transform (STFT) is computed with a window size equal to $2048$ and hop size equal to $512$ samples. This metric emphasizes the ability to predict portions of the output sound with rapid changes.
To account for the frequency response, we use the multi-resolution STFT  (Engel et al.,, 2020):
$\begin{split}M_{STFT}&=\frac{1}{N}\sum_{m}|||STFT_{m}(y_{n})|-|STFT_{m}(\hat{y}_{n})|||_{1}\\ &+\frac{1}{N}\sum_{m}||\log(|STFT_{m}(y_{n})|)-\log(|STFT_{m}(\hat{y}_{n})|)||_{1}\end{split}$ (10)
@@
which compares linear and logarithmic spectral distances with varying frequency resolutions using the L1 norm. The error is calculated as the sum of the absolute values of the vectors. This metric quantifies the error between the spectra employing different resolutions. We compute the metrics using $m=[256,512,1024]$ as resolutions.
Similarly, we compare the MFCCs to compute a metric linked to human sound perception:
$M_{MFCC}=\frac{1}{N}|||MFCC(y_{n})|-|MFCC(\hat{y}_{n})|||_{1}$ (11)
where $1024$ is the FFT length, and we have $80$ Mel bands between $30$ Hz and Nyquist frequency.

### v6

@@ 3 Methods
This study aims to identify which type of recurrent layer or architectural configuration accurately models different types of audio effects, if any. These insights can be used to understand the strengths and weaknesses of these models and determine which can better suit each audio effect. The state-based architectures offer advantages due to their ability to capture time dependencies in the data by utilizing internal states rather than solely relying on the input received at each iteration. In addition, we restrict the experimentation to relatively small networks that are fed by a limited number of input samples. This approach helps reduce computational complexity and minimizes input-output latency, enabling implementation on real-time consumer-grade digital audio systems.
The designed models are applied to various audio effects and evaluated using a range of metrics. It is important to note that, given their limited number of trainable parameters, these models are not expected to achieve perfect emulation but rather serve as a solid foundation for further development.
In the rest of this section, we first detail the five architectures used in this comparative study; then, we present the metrics we used for evaluating the trained models and the specific audio effects from which we collected the datasets. Finally, explain the process we used to train and test the models.
@@ 3.1 Architectures
The following architectures, based on different types of recurrent layers, are compared in this study: LSTM, ED, LRU, S4D, and S6. The five architectures are illustrated in Figure 1. ED is based on the concept of internal state sharing  Simionato & Fasciani, (2022). These architectures follow the same design, consisting of a linear, fully connected layer, a recurrent layer, and a conditioning block.
The number of input samples is identical for all architectures and set to $64$. This choice represents a trade-off: on the one hand, there is the amount of past input information used to generate the current output, where more is typically better; on the other hand, there is the resulting input-output latency, which increases with the number of input samples. Our goal is to keep this latency below $1.33$ ms (i.e., $64$ samples at $48$ kHz). The input of all architectures is an array, including the $64$ most recent input sample, used to generate the single current output sample. This implies that a new input sample enters the buffer with each inference cycle while the oldest is discarded. Therefore, the set of the $64$ most recent input samples constitutes the network input, which, after initial compression by the FC layer, is used to update the states in the recurrent layers. This approach aids the network in making predictions while maintaining a stateful design using truncated backpropagation through time. Predicting one sample at a time helps minimize the audible artifacts commonly associated with machine-learning models of audio effects, which arise from slight amplitude mismatches at the boundaries of consecutive output segments. Consequently, the model must perform predictions at an audio rate, one for each output audio sample. The models map a vector of $64$ samples into one scalar output sample, giving the internal states $\boldsymbol{h}$:
@@
where $\boldsymbol{u_{n}}$ is the linearly projected vector having a dimensionality equal to the number of hidden units of the fully connected layer. $\boldsymbol{W_{l}}$ and is the weights’ matrices, and $\boldsymbol{b_{l}}$ the vector bias terms.
Figure 1: The five compared architectures: LSTM, ED, LRU, S4D, and S6. All architectures present the same structure: a fully connected layer, a recurrent layer, and a conditioning layer. For the case of ED, the input vector is equally split into two vectors, $\boldsymbol{x_{e}}$ and $\boldsymbol{x_{d}}$. The vector $\boldsymbol{p}$ denotes the effect parameters utilized to condition the networks.
Following the linear projections, the models are characterized by various sets of equations, contingent upon the specific type of layer employed.
LSTM layer
The vector $\boldsymbol{u_{n}}$ is processed using a LSTM as the recurrent layer. LSTMs were designed to improve the memory capacity of the vanilla RNN by incorporating internal mechanisms known as gates, which regulate the flow of information. The LSTM architecture is governed by the following equations:
@@
where $\boldsymbol{h_{n}}$ is the hidden state vector, $\boldsymbol{f}$ is the forget gate, $\boldsymbol{i}$ the input gate, and $\boldsymbol{o}$ the output gate, $\boldsymbol{c}$ is the cell state vector, and $\circ$ the element-wise product. The forget gate determines what information from previous steps should be kept or discarded. The input gate decides which information from the current step is relevant to add. The output gate determines the value of the next hidden state. Additionally, the cell state is updated by multiplying the forget vector and the input vector with new values that the neural network deems relevant. The forget vector and input vector are obtained using the sigmoid function, denoted by $\gamma$, which determines how much information should be kept. Values of 0 in this vector indicate information to be discarded, while a value of $1$ indicates information to be kept. Furthermore, the $\phi$ represents the hyperbolic tangent function, used to keep the value of the cell state between 1-1 and $1$. Finally, $\boldsymbol{W}$ and $\boldsymbol{U}$ are the weights’ matrices, and $\boldsymbol{b}$ the vector bias terms.
ED layer
In this case, the input of the network $\boldsymbol{x}$ is split into two equal-size parts, $\boldsymbol{x_{d}}$ containing $x_{n},...,x_{n-31}]$ and $\boldsymbol{x_{e}}$ containing containing $x_{n-32},...,x_{n-63}]$. The first part feeds two separate convolutional layers, which compute the states, $\boldsymbol{h},\boldsymbol{c}]$. These states are combined with the internal states of the LSTM layer computed during the previous iteration. A sigmoid function is applied to the LSTM internal states and then element-wise multiplied by $\boldsymbol{h},\boldsymbol{c}]$. This operation determines how much information should be taken from the computed $\boldsymbol{h},\boldsymbol{c}]$. The result of this operation represents the new LSTM’s internal states. This process is detailed in Figure 2.
Figure 2: The states sharing process in ED: the encoder input $\boldsymbol{x_{e}}$ is passed to two convolutional layers, which produce the states $\boldsymbol{h},\boldsymbol{c}]$. These states are combined with $\boldsymbol{h_{n-1}},\boldsymbol{c_{n-1}}]$, which are the LSTM states of the previous iteration. The sigmoid function is applied to $\boldsymbol{h_{n-1}},\boldsymbol{c_{n-1}}]$ and the results are element-wise multiplied to $\boldsymbol{h},\boldsymbol{c}]$. The outcome will be the new states for the LSTM layer $\boldsymbol{h_{n}},\boldsymbol{c_{n}}]$.
LRU layer
Compared to a vanilla RNN, the LRU uses a linear activation function to compute the hidden state instead of a nonlinear function. This modification allows the memory capability to improve significantly. The LRU equations are the following:
@@
where $\boldsymbol{h_{n}}$ is the hidden state at time $n$ and acts as ’memory’ of the network, $\boldsymbol{o_{n}}$ the output vector. LRU keeps the recurrent layer linear but moves the nonlinearity to a subsequent fully connected layer. Additionally, to further enhance stability, the LRU incorporates a complex-valued diagonal recurrent matrix initialization and an exponential parameterization. At the same time, a normalization scheme for the hidden states is used to boost the efficiency and accuracy Orvieto et al., (2023).
S4D layer
The SSMs operate by considering a continuous-time representation of the state-space formulation that is discretized to obtain the following equations:
@@
In our study, we employ the diagonal state-space model variant S4D  Gu et al., (2022a). The remarkable ability of S4D to capture long-range dependencies stems from its utilization of a specific state matrix $\boldsymbol{\bar{W_{A}}}$ known as the "HiPPO matrix" Gu et al., (2020). This matrix is designed to encode the entire history of past inputs within $\boldsymbol{h_{n}}$, effectively mapping $\boldsymbol{u_{n}}$ to a higher-dimensional space $\boldsymbol{h_{n}}$ that encapsulates the compressed representation of this historical input data. The matrix allows the model to be conceptualized as a convolutional model that decomposes an input signal onto an orthogonal system of smooth basis functions. In this way, the state $\boldsymbol{h_{n}}$ encodes the history of the inputs. In the diagonal form, the S4D model is defined by parameterizing its state matrix as a diagonal matrix, boosting the model’s efficiency. By doing so, the basis kernels have closed-form formulas represented by normalized Legendre polynomials $L_{n}(t)$, resulting in the SSM decomposing the input $\boldsymbol{u_{n}}$ onto an infinitely long set of basis functions that are orthogonal with respect to an exponentially decaying measure.
S6 layer
The S6 layer Gu & Dao, (2023) introduces input-dependent matrices. Specifically, $\boldsymbol{\bar{W_{B}}}$ and $\boldsymbol{\bar{W_{C}}}$ become dependent on the input and computed using a linear FC layer, while $\boldsymbol{\bar{W_{A}}}$ and $\boldsymbol{\bar{W_{D}}}$’s parameters are independent of the input and learned during the training process.
After the recurrent layer—whether LSTM, ED, LRU, S4D or S6—a FC layer reduces the dimensionality to $4$, ensuring uniformity across all model variations before the conditioning block. For the LRU, S4D, and S6, this FC layer includes a hyperbolic tangent activation function, compensating for the absence of nonlinear activation functions within these layers. Conversely, the LSTM model uses a linear activation function in the FC layer, as detailed in the following equations:
$\boldsymbol{\hat{o}_{n}}$ $=tanh(\boldsymbol{W_{nl}}\boldsymbol{o_{n}}+\boldsymbol{b_{nl}})\text{ if }\textbf{LRU / S4D / S6}$
$\boldsymbol{\hat{o}_{n}}$ $=(\boldsymbol{W_{nl}}\boldsymbol{o_{n}}+\boldsymbol{b_{nl}})\text{ if }\textbf{LSTM}$ (6)
The conditioning block is then applied just before the output layer, which consists of a linear fully connected layer with one unit, which is described in the following subsection. Placing the conditioning block after the recurrent layers rather than before has been found beneficial  Fasciani et al., (2024a). This suggests that it is more beneficial for the networks to use the information given by the control parameters to project the output of the recurrent layer and determine the extent to which this information influences the final output rather than influencing the inference of the recurrent layer based on the control parameters.
@@
The number of units is chosen to have approximately $800$ trainable parameters among the architectures and effects, since an identical number is achievable due to major differences in the layers. As a result, the selected LSTM and ED present $8$ units in the recurrent layer and $4$ in fully connected ones. LRU, S4D and S6 present $12$ units in the recurrent layer, and $6$ in fully connected ones. The number of trainable parameters also depends on the dimensionality of the conditioning input $\boldsymbol{p}$, equal to the number of variable parameters in the datasets.
@@ 3.1.1 Conditioning Block
The conditioning block consists of the FiLM method and the GLU layer, as in Fasciani et al., (2024a), and is illustrated in Figure 3.
@@
FiLM is a technique that applies an affine transformation to a vector based on conditioning values. The parameters vector $\boldsymbol{p}$, if applicable, is the input to a linear FC layer. The output of this layer is then split into two vectors of equal size: $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$. If $\boldsymbol{o_{n}}$ is the vector to which the transformation is applied, the output of the linear FC layer is twice the length of $\boldsymbol{o_{n}}$. Consequently, FiLM performs an affine transform on $\boldsymbol{o_{n}}$ using $\boldsymbol{\theta}$ and $\boldsymbol{\eta}$:
$\boldsymbol{q}$ $=\boldsymbol{\theta}\boldsymbol{o_{n}}+\boldsymbol{\eta}.$ (8)
Following the projection, it is common to apply a function, such as ReLU or sigmoid, to determine the amount of information that should be passed. In our approach, we replace the activation function with a GLU layer. This layer is more stable than ReLU and exhibits faster learning compared to sigmoid. Similarly to the previous step, the GLU layer consists of a linear FC layer that takes the FiLM output vector as input and computes a vector twice its length. The resulting output is then divided equally into two vectors: $\boldsymbol{q_{1}}$ and $\boldsymbol{q_{2}}$. A function is applied to $\boldsymbol{q_{2}}$, with the resulting output being multiplied element-wise with $\boldsymbol{q_{1}}$. This operation describes the conditioning block, represented by the following equation:
$\boldsymbol{o^{c}_{n}}=\boldsymbol{q_{1}}\odot softsign(\boldsymbol{q_{2}})$ (9)
The GLU layer determines the flow of information through the network, acting as a logical gate. Typically, a sigmoid activation function is used for this purpose. However, in our approach, we introduce a softsign function instead. The softsign function controls the extent to which the control parameters should positively or negatively influence the final output. A value of 0 indicates a bypass, meaning no influence on the output. This allows for more flexibility.
@@ 3.2 Losses and Metrics
The models in this study take raw audio samples as input and produce raw audio samples as output. The most common losses used in this case include the Mean Squared Error (MSE), the Mean Absolute Error (MAE), and the Error-to-Signal Ratio (ESR). Based on our knowledge and supported by empirical experiments, there is no evidence suggesting that more complex loss functions could benefit the learning process, especially when modeling the transformation applied to a signal by a generic audio effect. While using specific and different loss functions for different effects may potentially yield better results, we decided to carry out this study using a fixed loss function across architectures and audio effects. This allows for a fair comparison between the trained models, as they all attempt to minimize the same loss function. Therefore, we selected the MSE for this comparative study as it helps to capture the dynamic aspects of the output signal more effectively  Simionato & Fasciani, (2023).
@@
In addition, our quantitative evaluation is based on various metrics to assess the models from different perspectives. In the following expressions, $\boldsymbol{y}$ and $\hat{\boldsymbol{y}}$ represent the target and predicted audio signals, and the symbol $N$ represents the overall number of samples in the output signal.
The first metric we utilize is the normalized root-mean-square energy (NRMSE) error, with normalization carried out using the RMSE of the target signal. NRMSE error provides insights into the energy deviation between the target and the prediction.
@@
The losses and metrics presented so far calculate the average difference between the target and prediction, which may hide large errors that occur only during signal transients. Transients, such as sharp attacks or decays, may represent a small fraction of the overall duration of an audio signal. In this case, we may obtain a relatively low error even with all transients poorly modeled. Therefore, we have included a metric related to the spectral flux, which considers the differences between pairs of spectra computed on consecutive overlapping windows, as used in the library  [^1] https://github.com/magenta/ddsp developed after  Engel et al., (2020):
[…]
where the Short-Time Fourier Transform (STFT) is computed with a window size equal to $2048$ and hop size equal to $512$ samples. The metric is normalized by the squared difference between the magnitudes of the spectra of two successive target frames, measuring the spectral change between them. This indicates the rate at which the power spectrum of a signal is evolving over time.
To account for the frequency response, we use the multi-resolution STFT  Engel et al., (2020):
[…]
@@
which compares linear and logarithmic spectral distances with varying frequency resolutions using the L1 norm. The linear component of $M_{STFT}$ is normalized by the magnitude of the STFT of the target. The error is calculated as the sum of the absolute values of the vectors. This metric quantifies the error between the spectra employing different resolutions. We compute the metrics using $m=[256,512,1024]$ as resolutions.